Solve the following simultaneous equations:
xy = x + 6
x – 2y = 2

Answer by anantha(86)

sol:xy=x+6 this is 1st equation
x-2y=2 this is 2nd equation
take 2nd equation
x-2y=2
adding 2y on both sides
x-2y+2y=2+2y
x=2+2y
substitute x value in 1st equation
xy=x+6
(2+2y)*y=2+2y+6
2*y+2*2y=8+2y
2y+4y=8+2y
6y=8+2y
subtracting 2y on both sides
6y-2y=8+2y-2y
4y=8
dividing 4 on both sides
4y/4=8/4
y=2
substitute this value in 2nd equation we get x value
x-2y=2
x-2*2=2
x-4=2
adding 4 on both sides
x-4+4=2+4
x=6
check these values satisfie the equation or not
take 2nd equation
x-2y=2
6-2*2=2
6-4=2
2=2
take 1st equation xy=x+6
6*2=6+6
12=12
these values satisfy both equations